OpenStudy (anonymous):

A triangle has the sides 20 cm and angles 15 degrees and 125 degrees. 1. measure of the other angle? is it 40? 2. measure of other sides? are they 6.3 and 15.7?

OpenStudy (anonymous):

I used the law of sines.

OpenStudy (mertsj):

Can't tell unless we know the relationship between the given side and the angles.

OpenStudy (anonymous):

In triangle ABC (angles), A is unknown, C is 125, and B is 15. The side length c (across from angle C is 20

OpenStudy (anonymous):

i got approximately 6.3 for side b

OpenStudy (mertsj):

\[\frac{\sin125}{20}=\frac{\sin15}{b}, b=6.3\]

OpenStudy (anonymous):

okay thats what I did, is angle A forty and side a 15.6?

OpenStudy (anonymous):

approx 15.7 for side c

OpenStudy (mertsj):

\[\frac{\sin125}{20}=\frac{\sin40}{a}, a=15.7\]

OpenStudy (anonymous):

okay thank you! and if i use herons formula to find the area, is it 40.6?

OpenStudy (anonymous):

yes, a, my mistake